OpenStudy (anonymous):

How would I set this up? Add and simplify: (10 x)/3 + (-x)/3

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

help me guys

OpenStudy (anonymous):

Well the result would be 3 x. But what do you mean how do you set it up?

OpenStudy (anonymous):

Because I found the answer to be 3x as well and it says my answer is wrong so I thought I was not doing it properly.

OpenStudy (shamil98):

\[\frac{ 10x }{ 3 } - \frac{ x }{ 3 } = \frac{ 9x }{ 3 }=\frac{ \cancel 9 x }{ \cancel 3 } = 3x\]